OpenStudy (anonymous):

3/4 divided by 1/9 (write your answers in lowest terms and as whole or mixed number if possible)

hero (hero):

\[\frac{3}{4} \div\ \frac{1}{9} \\= \frac{3}{4} \times \frac{9}{1} \\= \frac{27}{4} \\= \frac{28 - 1}{4} \\= \frac{28}{4} - \frac{1}{4} \\= 7 - \frac{1}{4} \\= 6 + \frac{4}{4} - \frac{1}{4} \\= 6 + \frac{3}{4}\]